Find three consecutive numbers such that the sum of the first and twice the second is forty-six less than six times the third. The third number is
Answer by ewatrrr(24785)

x + 2(x+1)= 6(x+2)-46***
x + 2x + 2 = 6x + 12 - 46
36 = 3x
12 = x
three consecutive numbers are: 12, 13, 14
checking***: 12+26 = 6*14 -46 = 38 Checks